Three charges 4q,Q and q are in straight line in the position o,l/2 and l respectively the resultant force on q will be  zero

Grade:12

3 Answers

Badiuddin askIITians.ismu Expert
14 years ago

Dear pallavi

two force are acting on q ,one due to 4q and second due to Q

F1 = 1/4∏ε (4q)q/l

F2 = 1/4∏ε (Q)q/(l/2)

      = 1/4∏ε (4Q)q/l

 F1 +F2 =0

1/4∏ε (4q)q/l2   +  1/4∏ε (4Q)q/l=0

 Q=-q

 Please feel free to post as many doubts on our discussion forum as you can.
 If you find any question Difficult to understand - post it here and we will get you the answer and detailed solution very quickly.
 We are all IITians and here to help you in your IIT JEE  & AIEEE preparation.

 All the best.
 
Regards,
Askiitians Experts
Badiuddin

Kalai
5 years ago
three point charges 4q q and 2q are placed in a straight line of length 10 cm at points distance 0.5 cm 10 cm respectively.if the net force on charge 2q is zero,then calculate the value of q
Rishi Sharma
3 years ago
Dear pallavi
two force are acting on q ,one due to 4q and second due to Q
F1 = K(4q)q/1^2
F2 = K(Q)q/(1/2)^2
= K(4Q)q/1
F1 +F2 =0
K(4q)q/1^2 + K(4Q)q/1 =0
Q=-q
Thanks
I hope above solution solve your doubts.
Please feel free to post as many doubts on our discussion forum as you can.
All the best.